# Present address: ]Savannah River National Laboratory, Savannah River Site, SC, 29803 ]www.amberlauercoles.com

### v1

@@ VI Conclusions
Resonances of 37K in 38Ca have been studied via proton scattering with a 37K beam on a CH2 target at the ReA3 facility using a Si detector array and a position-sensitive ionization chamber to determine the excitation function between approximately MeV in the 38Ca compound nucleus. In addition to the identification of 13 new levels, the Jπ values and proton partial widths of these levels, as well as for levels newly identified by Long et al. Long et al. (2017) in this energy range, have been constrained through an R-Matrix analysis of these results. This information was used to estimate a new 34Ar()37K reaction rate, which was found to be a factor of 45 less than the standard REACLIB reaction rate, based on Hauser-Feschbach theory. While similar to other indirect studies Long et al. (2017), the rate determined here does not agree with the only previous higher-energy direct measurement Browne et al. (2023), which is itself consistent with HF theory.
The sensitivity to this new reaction rate of an XRB model based on GS 1826-24 in MESA astrophysical simulations was studied by comparing the outputs of a baseline model run with the current recommended reaction rate in REACLIB, as well as variations of that rate by factors of 100 up and down, and the rate determined in this work. The results showed no significant deviations in the luminosity profile, burst recurrence time, or burst duration of this initial model when the 34Ar()37K rate is varied; however, large variations in the maximum luminosity due to this rate were observed for altered models with slower accretion rates or a larger He fraction. While this does not agree with the results from the single zone models described in Cyburt et al. Cyburt et al. (2016) or Browne et al. Browne et al. (2023), these results are similar to those found in the multizone model of Cyburt et al.. It is possible these results could serve as caution against relying heavily on post-processed single-zone models, especially when the nuclear network includes no feedback with the thermodynamics, as is often the case. The results of the variation in abundances may indicate that a 300-species network is not large enough to faithfully explore the XRB scenario. This also suggests that the accepted results, both for this astrophysical scenario and others, derived from a decade or more ago, may need to be revisited considering the vast improvement in computational capabilities over the intervening years. While these results suggest a more precise determination of the 34Ar()37K reaction rate is not necessary to better understand XRB nucleosynthesis in GS 1826-24, such a measurement may be useful in determining the applicability of more exotic burst scenarios, as well as Hauser-Feschbach methods in this mass range given the discrepancies between these results and the accepted HF rate.
@@ VII Acknowledgments
This work was partially supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under contract numbers DE-FG02-96ER40978, DOE DE-FG02-02ER41220, DE-AC05-00OR22725, DE-FG02-93ER40773, and DE-AC02-06CH11357, Louisiana Board of Regents RCS Subprogram Contract LEQSF(2012-15)-RD-A-07, and the National Science Foundation under contract number PHY-1401574.

### v2

@@ VI Conclusions
Resonances of 37K in 38Ca have been studied via proton scattering with a 37K beam on a CH2 target at the ReA3 facility using a Si detector array and a position-sensitive ionization chamber to determine the excitation function between approximately MeV in the 38Ca compound nucleus. In addition to the identification of 13 new levels, the Jπ values and proton partial widths of these levels, as well as for levels previously identified by Long et al. Long et al. (2017) in this energy range, have been constrained through an R-Matrix analysis of these results. This information was used to estimate the 34Ar()37K reaction rate, which was found to be approximately a factor of 20-40 less than the statistical model rates that have been shown to describe the cross section at higher energies.
The sensitivity of XRB models to the 34Ar()37K rate was studied using MESA astrophysical simulations based on GS 1826-24. Outputs of a baseline model run with the current recommended reaction rate in REACLIB were compared to those with variations of that rate by factors of 100 up and down, as well as the rate determined in this work. We find no significant deviations in the luminosity profile, burst recurrence time, or burst duration of this model when the 34Ar()37K rate is varied; however, large variations in the maximum luminosity due to this rate were observed for models with slower accretion rates or a larger He fraction. This does not agree with the results from the single zone models described in Cyburt et al. Cyburt et al. (2016) or Browne et al. Browne et al. (2023), nor with the results found in the KEPLER multizone model of Cyburt et al. when the 34Ar()37K rate was varied by the same factor of 100 as in this work. Such differences between multizone and single zone models have been observed previously Cyburt et al. (2016) and are not surprising, while the differences between the MESA models presented here and the multizone models of Cyburt et al. may be due to the fact that this reaction is near the endpoint of the process and as such small differences in the temperature profile as a function of depth may have large effects on the models. It is possible these results could serve as caution against relying heavily on post-processed single-zone models, especially when the nuclear network includes no feedback with the thermodynamics, as is often the case (although Cyburt et al. (2016) did include feedback in their one zone model). This also suggests that the sensitivity of various models to different reactions can vary based on the model’s parameters, architecture, etc., especially given the vast improvement in computational capabilities over the intervening years.
@@ VII Acknowledgments
This work was partially supported by the U.S. Department of Energy, Office of Science contract no. DE-SC0014231 and Office of Nuclear Physics under contract numbers DE-FG02-96ER40978, DOE DE-FG02-02ER41220, DE-AC05-00OR22725, DE-FG02-93ER40773, and DE-AC02-06CH11357, Louisiana Board of Regents RCS Subprogram Contract LEQSF(2012-15)-RD-A-07, the National Science Foundation under contract number PHY-1401574, and IBS grant number IBS-R031-D1, funded by the Korea government.